If a 180 - mile trip took 71/2 gallons of gas ,how many miles can be driven with a full tank of 14 gallons?

--mi
1
Expert's answer
2013-02-04T09:25:30-0500
If a 180 - mile trip took 71/2 gallons of gas, then the consumption of gas is

V = 71/2[g]/180[m] = 71/360[g/m].

Therefore,

M = 14[g]/(71/360)[g/m] = 5040/71[m] ≈ 70.986[m]

miles can be driven with a full tank of 14 gallons.